Calculate the standard deviation and mean deviation from mean if the frequency function f(x) has the form: f(x)={3+2x/18,for2≤x≤4; 0 otherwise
1
Expert's answer
2020-03-21T17:24:15-0400

The frequency function f(x) has the form:


"f(x) = \begin{cases}
 {3+2x \over 18}, &\text{for } 2\leq x\leq4 \\
 0, &\text{otherwise }
\end{cases}"

"\displaystyle\int_{-\infin}^\infin f(x)dx=\displaystyle\int_{2}^4 {3+2x \over 18}dx=\bigg[{3x+x^2 \over 18}\bigg]\begin{matrix}
 4 \\
 2
\end{matrix}=""={3(4)+(4)^2 \over 18}-{3(2)+(2)^2 \over 18}=1"

"\mu=E(X)=\displaystyle\int_{-\infin}^\infin xf(x)dx=\displaystyle\int_{2}^4 x{3+2x \over 18}dx="

"=\bigg[{{3x^2 \over 2}+{2x^3 \over 3} \over 18}\bigg]\begin{matrix}
 4 \\
 2
\end{matrix}={9(4)^2+4(4)^3 \over 108}-{9(2)^2+4(2)^3 \over 108}={83 \over 27}"

"E(X^2)=\displaystyle\int_{-\infin}^\infin x^2f(x)dx=\displaystyle\int_{2}^4 x^2{3+2x \over 18}dx="

"=\bigg[{{3x^3 \over 3}+{2x^4 \over 4} \over 18}\bigg]\begin{matrix}
 4 \\
 2
\end{matrix}={2(4)^3+(4)^4 \over 36}-{2(2)^3+(2)^4 \over 36}={88 \over 9}"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2="

"={88 \over 9}-\big({83 \over 27}\big)^2={239 \over 729}"


"\mu=E(X)={83 \over 27}"

"Var(X)=\sigma^2={239 \over 729}"

"\sigma_X=\sqrt{Var(X)}=\sqrt{{239 \over 729}}\approx0.5726"

"MD=\displaystyle\int_{-\infin}^\infin (x-\mu)f(x)dx="

"=\displaystyle\int_{2}^4 (x-{83 \over 27}){3+2x \over 18}dx="

"=\bigg[{36x^3-85x^2-498x\over 972}\bigg]\begin{matrix}
 4 \\
 2
\end{matrix}=0"